OpenStudy (anonymous):

Combine like terms to simplify the expression: 2/5m-4/5-3/5m

Nnesha (nnesha):

\[\huge\rm \frac{ 2 }{ 5 }m-\frac{ 4 }{ 5 }-\frac{ 3 }{ 5 }m\] ? like this

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }m-\frac{ 4 }{ 5 }-\frac{ 3 }{ 5 }m\]

Nnesha (nnesha):

alright just add the coefficient of same variable here are some example \[\huge\rm 3\color{reD}{x}+2\color{ReD}{x} \] variables are same so you can combine their numbers (3+2)x=5x

OpenStudy (anonymous):

im so lost just instruct me though mine ill get sidetracked and lost on what im doing plz

Nnesha (nnesha):

okay that is just an example of *like terms * in your question \[\huge\rm \color{reD}{\frac{ 2 }{ 5 }m}-\frac{ 4 }{ 5 }-\color{reD}{\frac{ 3 }{ 5 }m}\] red ones are like terms bec they both have the same variable combine them \[(\frac{ 2 }{ 5 }m-\frac{ 3 }{ 5 }m)-\frac{ 4 }{ 5}\]

OpenStudy (anonymous):

ok got tht

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

alright that's it!

Nnesha (nnesha):

there are just 2/5 , -3/5 like terms

OpenStudy (anonymous):

uggh nevermint i had a bug so soerry to waste ur time

Nnesha (nnesha):

it's okay :=)